# RecGS: Removing Water Caustic with Recurrent Gaussian Splatting

### v1

@@ Abstract
Water caustics are commonly observed in seafloor imaging data from shallow-water areas. Traditional methods that remove caustic patterns from images often rely on 2D filtering or pre-training on an annotated dataset, hindering the performance when generalizing to real-world seafloor data with 3D structures. In this paper, we present a novel method Recurrent Gaussian Splatting, which takes advantage of today’s photorealistic 3D reconstruction technology, 3DGS (3DGS), to separate caustics from seafloor imagery. With a sequence of images taken by an underwater robot, we build 3DGS recursively and decompose the caustic with low-pass filtering in each iteration. In the experiments, we analyze and compare with different methods, including joint optimization, 2D filtering, and deep learning approaches. The results show that our method can effectively separate the caustic from the seafloor, improving the visual appearance. Our videos and code are released at https://https://tyz1030.github.io/proj/recgs.html
@@ I INTRODUCTION
Within the literature of 3D reconstruction, SfM (SfM) methods are some of the most fundamental. These methods recover the camera poses and sparse 3D structure of the scene from a sequence of images [buildrome]. However, the visualization results of SfM are generally sparse and not photorealistic. Therefore, recent advancements in 3D reconstruction such as NeRF (NeRF[mildenhall2020nerf] and 3DGS [kerbl3Dgaussians] build on top of the results of SfM to enable photorealistic novel view rendering of the scene. These methods constitute the SOTA (SOTA) of 3D reconstruction. In general, SfM pipelines are robust against illumination change due to sophisticated outlier rejection mechanisms. Therefore, they can be applied to internet data with different levels of ambient lighting and camera parameters [buildrome], while still capable of accurately estimating the poses of the cameras. We also observe that water caustics generally have a limited impact on traditional SfM pipelines for the same reason. Recovering photorealistic visualisations is of particular significance underwater and has implications for other branches of science. However, water caustics are detrimental to the construction of photorealistic representations, which are built on top of SfM. These include NeRF and 3DGS which further learn the 3D scene by optimizing photometric loss instead of re-projection error. As a result, these models are critically affected by the illumination inconsistency from caustics.
@@
Existing studies on caustic removal are based either on image filtering or supervised deep learning. Filtering-based methods [flicker04, filter2008oceans, flicker12] estimate and separate the caustic by running a non-linear temporal filter with motion compensation. However, such hand-crafted filtering techniques assume that the motion is small and has no awareness of 3D structure, thus the performance can be easily downgraded due to non-optimal hyperparameters, increased motion speed, and seafloor rugosity. Deep learning methods [deepcaustics2019, seafloorinv2023] require an annotated dataset on which to train. However, such dense caustic annotations are extremely expensive to acquire, hard to scale up, and only monochromatic [seafloorinv2023]. Therefore, methods based on deep learning perform poorly when generalizing to novel observations.
In this paper, we study caustic removal as a problem of illumination inconsistency within 3D structures. The community has seen related problems with color [neuralsea], brightness [zhang2024darkgs] and exposure inconsistency [darmon2024robust, wang2024bilateral] being addressed with 3D vision techniques without any annotations. We seek to push the boundaries of this line of work and exploit the 3D consistency maintained in representation to remove caustics. To this end, we present recurrent Gaussian Splatting, a novel framework which applies iterative low-pass filtering to refine the quality of the 3D representation during construction. Compared with existing approaches for water caustic removal, our proposed approach is unique in its structure-awareness and unsupervised nature. Namely, our approach leverages the consistency within the 3D structure of our representation to enhance performance and does not demand any annotated labels, enabling efficient scaling-up to large scenes. We provide extensive empirical evaluations on our underwater dataset, of shallow water corals, collected with a ROV (ROV) in Florida Keys, USA. Our method consistently and robustly produces clean photorealistic representations of underwater scenes that are caustic-free. A high-level illustration of our problem setup and method is provided in Fig. 1.
@@ IV EXPERIMENTS
In this section, we provide rigorous empirical evaluations of our proposed recurrent Gaussian Splatting method. We provide details of our experiment setup in section IV-A. Then, seek to investigate and provide answers to the following questions:
  • Should the caustic effects be decomposed or modelled jointly? (Section IV-B)
@@ IV-C Failure case of Pre-trained Deep Learning Methods
(b) Undesired caustics due to improperly selected hyper-parameter.
Figure 9: Limitations of our method.
The scaling law has played a critical role on a variety of computer vision tasks in the past decade, that with larger amount of data and model size, the model performance gets significantly improved. Deep learning approaches to caustic removal seek to train on a sufficiently large dataset with labelled data and generalise when provided images, under potentially different conditions and at new locations. However, the community has seen much less progress in scaling up underwater computer vision studies, as the real-world data are extremely expensive to collect and need expert input to label. The SOTA data for learning caustic is RCaustic [seafloorinv2023], which contains 712 pairs/triplets of images from 7 scenes with annotated caustic patterns and contours. However, for the monocular camera setup, neural networks trained on such a dataset perform poorly when transferring to novel data collected from the field. In Fig. 6, we show that the visual appearance gap between the training data and real-world data, and display an example of the constant failure of pre-trained neural networks when domain gap presents. Our proposed recurrent Gaussian Splatting approach does not suffer the same drawbacks, as it does not rely on training on a labelled dataset to generalize to novel images.

### v2

@@ Abstract
Water caustics are commonly observed in seafloor imaging data from shallow-water areas. Traditional methods that remove caustic patterns from images often rely on 2D filtering or pre-training on an annotated dataset, hindering the performance when generalizing to real-world seafloor data with 3D structures. In this paper, we present a novel method Recurrent Gaussian Splatting (RecGS), which takes advantage of today’s photorealistic 3D reconstruction technology, 3DGS (3DGS), to separate caustics from seafloor imagery. With a sequence of images taken by an underwater robot, we build 3DGS recurrently and decompose the caustic with low-pass filtering in each iteration. In the experiments, we analyze and compare with different methods, including joint optimization, 2D filtering, and deep learning approaches. The results show that our proposed RecGS paradigm can effectively separate the caustic from the seafloor, improving the visual appearance, and can be potentially applied on more problems with inconsistent illumination. Our videos and code are released at https://tyz1030.github.io/proj/recgs.html
@@ I INTRODUCTION
Within the literature of 3D reconstruction, SfM (SfM) methods are some of the most fundamental. These methods recover the camera poses and sparse 3D structure of the scene from a sequence of images [buildrome]. However, the visualization results of SfM are generally sparse and not photorealistic. Therefore, recent advancements in 3D reconstruction such as NeRF (NeRF[mildenhall2020nerf] and 3DGS [kerbl3Dgaussians] build on top of the results of SfM to enable photorealistic novel-view rendering of the scene. These methods constitute the SOTA (SOTA) of 3D reconstruction. In general, SfM pipelines are robust against illumination change due to sophisticated outlier rejection mechanisms. Therefore, they can be applied to internet data with different levels of ambient lighting and camera parameters [buildrome], while still capable of accurately estimating the poses of the cameras. We also observe that water caustics generally have a limited impact on traditional SfM pipelines for the same reason. Recovering photorealistic visualizations is of particular significance underwater and has implications for other branches of science. However, water caustics are detrimental to the construction of photorealistic representations, which are built on top of SfM. These include NeRF and 3DGS which further learn the 3D scene by optimizing photometric loss instead of re-projection error. As a result, these models are critically affected by the illumination inconsistency from caustics.
@@
Existing studies on caustic removal are based either on image filtering or supervised deep learning. Filtering-based methods [flicker04, filter2008oceans, flicker12] estimate and separate the caustic by running a non-linear temporal filter with motion compensation. However, such hand-crafted filtering techniques assume that the motion is small and has no awareness of 3D structure, thus the performance can be easily downgraded due to non-optimal hyperparameters, increased motion speed, and sealoor rugosity. Deep learning methods [deepcaustics2019, seafloorinv2023] require an annotated dataset on which to train. However, such dense caustic annotations are extremely expensive to acquire, hard to scale up, and only monochromatic [seafloorinv2023]. Therefore, methods based on deep learning perform poorly when generalizing to novel observations.
In this paper, we study caustic removal as a problem of illumination inconsistency within 3D structures. The community has seen related problems with color [neuralsea], brightness [zhang2024darkgs] and exposure inconsistency [darmon2024robust, wang2024bilateral] being addressed with 3D vision techniques without any annotations. We seek to push the boundaries of this line of work and exploit the 3D consistency maintained in representation to remove caustics. To this end, we present Recurrent Gaussian Splatting (RecGS), a novel framework which applies iterative low-pass filtering to refine the quality of the 3D representation during construction. Compared with existing approaches for water caustic removal, our proposed approach is unique in its structure-awareness and unsupervised nature. Namely, our approach leverages the consistency within the 3D structure of our representation to enhance performance and does not demand any annotated labels, enabling efficient scaling-up to large scenes. We provide extensive empirical evaluations on our underwater dataset, of shallow water corals, collected with a ROV (ROV) in Florida Keys, USA. Our method consistently and robustly produces clean photorealistic representations of underwater scenes that are caustic-free. A high-level illustration of our problem setup and method is provided in Fig. 1.
@@ IV EXPERIMENTS
In this section, we provide rigorous empirical evaluations of our proposed RecGS method. We provide details of our experiment setup in section IV-A. Then, seek to investigate and provide answers to the following questions:
  • Should the caustic effects be learned and optimized jointly with 3DGS? (Section IV-B)
@@ IV-C Failure case of Pre-trained Deep Learning Methods
The scaling law has played a critical role on a variety of computer vision tasks in the past decade, that with larger amount of data and model size, the model performance gets significantly improved. Deep learning approaches to caustic removal seek to train on a sufficiently large dataset with labelled data and generalise when provided images, under potentially different conditions and at new locations. However, the community has seen much less progress in scaling up underwater computer vision studies, as the real-world data are extremely expensive to collect and need expert input to label. The SOTA data for learning caustic is RCaustic [seafloorinv2023], which contains 712 pairs/triplets of images from 7 scenes with annotated caustic patterns and contours. However, for the monocular camera setup, neural networks trained on such a dataset perform poorly when transferring to novel data collected from the field. In Fig. 6, we show that the visual appearance gap between the training data and real-world data, and display an example of the constant failure of pre-trained neural networks when domain gap presents. Our proposed RecGS approach does not suffer the same drawbacks, as it does not rely on training on a labeled dataset to generalize to novel images.
@@ V LIMITATIONS
(b) Undesired caustics due to improperly selected hyper-parameter.
Figure 9: Limitations of our method.